OpenStudy (anonymous):

John began his job making $20 the first day. After that he was paid $6.00 per hour. Y = 6x + 20. Using the graph you drew for the equation, predict the pay for John if he works 7 hours after the first day.

OpenStudy (anonymous):

plug in 7 y = 6(7) + 20 or y = 42 + 20 can you get the last step?

OpenStudy (anonymous):

just add now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then 62 is my answer

OpenStudy (anonymous):

Thanks :D